# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v1

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing reasoning accuracy of Large Language Models (LLMs). However, current benchmarks mainly focus on closed tasks, leaving a gap in the assessment of more complex, real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both the reduction in hallucinations and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we further propose another experiment setting OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on methods and future directions for leveraging KGs to reduce LLMs’ hallucination. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs. The code of this paper is released at https://anonymous.4open.science/r/OKGQA-CBB0.
@@ 1 Introduction
To address this critical limitation, researchers employ diverse strategies to augment the LLMs by integrating external knowledge graphs (KGs) (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
@@
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific facts of KG. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios. Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
@@
Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
@@
To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types, and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities. Specifically, OKGQA covers 850 queries from ten categories, representing different levels of real-world QA types. All queries are synthesized by LLMs with pre-defined templates mentioned in Table 1, and we use Wikipedia as references for all the entities mentioned in the query. All the questions are designed to be open-ended and cannot be simply answered by generating logical forms or retrieving specific facts from the knowledge graph. To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (discussed in Section 3.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We assess the hallucination ratio and the quality of the response (detailed metrics can be found in Section 3.3) of the tested models.
@@
Based on our experiments, we find that (1) retrieving KG information can indeed mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree. Overall, our main contributions include:
  • We present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios. The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
  • To consider the scenarios in which KGs may have varying levels of mistakes, we propose another experiment setup OKGQA-P to assess model performance under conditions in which the semantics and structure of KGs are deliberately perturbed and contaminated.
  • We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for future research and development. Our results underscore the importance of integrating LLMs with KGs to help reduce hallucinations, even in circumstances where KGs are contaminated.
2 Related Work
Due to the stochastic decoding process of Large language models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behavior, potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023). In addition, they also face challenges in accurately interpreting phrases or terms when the context is vague and resides in a knowledge gap region of the model, leading to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations" undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs. For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval" to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, Reasoning on Graphs (RoG) (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable and accurate reasoning in LLMs.
@@ 3 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
OKGQA is a benchmark designed to assess LLMs enhanced with KGs under open-ended, real-world questions answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancement in reasoning capabilities. In this section, we discuss the construction of the dataset, and the construction of OKGQA-P which is a variant of OKGQA where the KGs’ semantics and structure are deliberately perturbed and contaminated.
Motivation of Open-ended QA instead of Close-ended QA: Current benchmarks for testing the capabilities of LLM+KGs models are predominantly close-ended [^1] https://en.wikipedia.org/wiki/Open-ended_question#Examples, which demand a short answer such as ‘yes’ and ‘no’. This restricts the responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific KG facts. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of more complex, real-world scenarios. Particularly, standard metrics (like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024)) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
Open-ended QA Task Definition: The tasks we design require to first understand the scope of the question, then optionally retrieve relevant information from multiple parts of the knowledge graph, then finally synthesize a coherent and informative response. The ideal output should be a paragraph that fully addresses the question with accurate and factual responses. We verify the response based on the metrics specified in Section 5 to reflect the models’ capabilities and faithfulness.
@@ 3.1 Dataset Construction
Queries. We utilize a template-based methodology to systematically generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding the specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix A.1. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to enhance the instructions for generating (the details are given in Appendix A.4). Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$. We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. These queries are categorized by difficulty and the naturalness of their phrasing. The statistics of the queries can be found in Figure 1.
@@
Table 1: Query types and examples in OKGQA. Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal refers to the specific entities to replace the placeholders. The distribution of the sub-query types can be referred to in Figure 1(b).
@@
KG Sub-graphs. To reduce the size of KGs, following previous works (Luo et al., 2023a; Yih et al., 2016; Talmor & Berant, 2018), we construct subgraphs of DBpedia (all the queries in OKGQA can be reasoned based on DBPedia [^2] https://www.dbpedia.org/resources/knowledge-graphs/) by extracting all triples contained within the k-hop neighbors from the question entities in the query. For our experiments, we set k=2. To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.5). We compare the statistics of subgraphs before and after PPR pruning in Figure 1(a).
3.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P in this section to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. Considering that KGs are typically annotated by humans and are generally accurate and meaningful, we introduce perturbations to edges in the KG to degrade the quality of the KGs, diminishing human comprehensibility. To quantify the degree of perturbation, we evaluate both the semantic and structual similarity between the original and the modified KG as defined below.
Notation. Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed from existing entities and relations. Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E^{\prime}},\mathcal{T^{\prime}})$ be the KG obtained after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$. Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$. Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs. Intuitively, ATS leverages an pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measures the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each of the three metrics, higher value indicates higher similarity. The detailed description can be found in Appendix A.7. The results after perturbation can be found in Figure 4.
For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations. Relation Replacement (RR) randomly chooses an edge $v_{1},e,v_{2})\in\mathcal{T}$, then replaces $e_{1}$ with another relation $e_{2}=\text{argmin}_{r\in\mathcal{R}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(v_{1},e,v_{2})$ uses ATS to measure the semantics similarity between two edges. Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$. Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it. We control the perturbation level based on the percentage of KG edges being perturbed.

### v2

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs). However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination. The code and data are released at link.
Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering
Yuan Sui1, Yufei He1, Zifeng Ding2, Bryan Hooi1 1 National University of Singapore 2 University of Cambridge {yuansui, yufei.he, bhooi}@comp.nus.edu.sg, zd320@cam.ac.uk
@@ 1 Introduction
Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
@@
To address this critical limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021). KGs provide several key advantages: they contain structured and explicit factual knowledge, including domain-specific information, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Furthermore, KGs offer traceability, allowing each piece of information to be traced back to its source. This traceability enables both verification of generated context and transparency in reasoning paths, making KGs a promising source for improving the reliability of LLM reasoning. Find more details of related works in §B.
@@
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities or relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor and Berant, 2018; Puerto et al., 2023) derived from specific facts of KG (as shown in Figure 1). These benchmarks, while useful for evaluating basic retrieval and reasoning, are insufficient for testing the LLMs’ ability for faithful generation (i.e., minimize hallucination in LLM responses), leaving a gap in the assessment of complex, real-world scenarios.
Why Open-ended Question? A key limitation of closed-ended benchmarks is: they cannot assess the phenomenon of “more talk, more mistakes”, where longer, more complex responses tend to increase hallucination Qiu et al. (2024). Unlike closed-ended benchmarks that require short, specific answers, open-ended questions demand more detailed and nuanced explanations. This format also enables direct evaluation metrics like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), which measure factual consistency of LLM responses by decomposing them into atomic statements for comparison with external information. However, for short, rigid responses, this decomposition is not feasible, rendering these metrics inapplicable. Moreover, open-ended evaluations align more closely with human cognition Lu et al. (2024); Amirizaniani et al. (2024), where context-rich and interpretative answers are expected compared to simplistic and mechanical responses. Find more details of intuition in Appendix §A.1.
@@
Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
@@
To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities. Specifically, all queries in OKGQA are synthesized by LLMs with pre-defined templates mentioned in Table 1 and corresponding Wikipedia context. To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We assess the hallucination ratio and the quality of the response (detailed metrics can be found in §4.1) of the tested models.
@@
Based on our experiments, we find that (1) retrieving KG information can mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
@@ 2 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
Overview. OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. The benchmark evaluates how effectively models leverage KGs to generate both reliable and comprehensive answers, focusing on reducing hallucinations while maintaining reasoning capabilities of LLMs. To assess robustness under imperfect conditions, we also introduce OKGQA-P, a variant where KGs are deliberately perturbed and contaminated.
2.1 Dataset Construction
@@
Queries. We utilize a template-based methodology to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix C. To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.2). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics. Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range. To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$. This optimization process ensures alignment between automated and human quality assessments. The queries are also categorized by complexity with detailed statistics in Figure 2.
KG Sub-graphs. To reduce the size of KGs, following previous works (Yih et al., 2016; Talmor and Berant, 2018), we sample subgraphs from DBpedia Knowledge Graph (Noted that all the queries in OKGQA can be reasoned based on DBPedia) by extracting all triples contained within the k-hop neighbors from the question entities in the query. We set k = 2 to balance graph size and computational feasibility, as increasing k leads to exponential growth in the number of edges and nodes, which can introduce excessive noise and make it more challenging to retrieve essential information from KGs [^1] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.3). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test the model’s reasoning capabilities with flawed or incomplete KG data. To quantify the degree of perturbation, we also evaluate both the semantic and structural similarity between the original and the modified KG as defined below.
Notation. Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed of entities and relations. Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ be the KG after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$. Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$. Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs. Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each of the three metrics, higher value indicates higher similarity. The detailed description can be found in Appendix A.5, with visualization in Figure 5.
For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations. Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\text{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS. The replacement is performed such that the resulting $(v_{1},e_{2},v_{2})$ has the closest score to the original triplet $(v_{1},e_{1},v_{2})$ to generate “harder negatives” - triplets that are semantically similar but incorrect, thereby challenging the model to different subtle distinctions. Noted that $S_{G}$ differs from the original ATS, as it operates on individual triplets rather than comparing two entire KGs. Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$. Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it. We control the perturbation level based on the percentage of KG edges perturbed. See results in Figures 5 and 6.
@@ 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
Table 1: Query types and examples in OKGQA. Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal highlights the specific entities to replace the placeholders.
@@
3 Exploring KG-augmented framework for reducing hallucination
To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigate a range of methods and discover that they can be encapsulated within a unified framework as shown in Figure 3. Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation. It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic design choices. We analyze our proposed strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
Figure 3: Overview of KG-augmented framework.
Here we formalize the problem of context-relevant KG retrieval for open-ended question answering. Given a user query $q$, a pretrained language model (PLM) first encodes the input tokens, and models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$. To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 3). This retrieved knowledge contains information relevant to the user query, and either benefits the LLMs at its knowledge boundaries (Baek et al., 2023b) or explicitly guides its reasoning process (Luo et al., 2023a; Sun et al., 2023). The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
where $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ represents the optimal subset of retrieved knowledge from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the likelihood generating the response given the query $q$ and the retrieved knowledge $\mathcal{Z}^{*}$, parameterized by $\phi$ and $\theta$. Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required. Consequently, the first line of Eq. 1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
3.1 Graph-guided retrieval (G-retrieval)
In this section, we explore which forms of knowledge (e.g., triplets, paths, or subgraphs) should be extracted from KGs to enhance the LLMs and discuss the design of G-retrieval. The phase begins by encoding the user query into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$. For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are similarly computed using the same model, ensuring all the information is represented in a unified vector space. The relevance of each node or edge to the query is measured using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$. Based on these similarity scores, the most relevant nodes and edges are identified to form a set of top-$k$ nodes and edges for the query.
To balance retrieving as many relevant nodes and edges as possible while keeping the size manageable, we adopt a prize-cost trade-off strategy: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs. These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fixed number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints. Finally, the top-scoring paths are sorted and returned. The details of path-retrieval can be found in Appendix A.4.
Sub-graph retrieval: the sub-graph retrieval approach also uses the prize assignment and cost allocation, following He et al. (2024a) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm. The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges. The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
3.2 Graph-guided generation (G-generator)
After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase learns to use this knowledge to generate the corresponding output to answer the user query. The generation is modeled as a sequential decision-making process, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)\vspace{-2mm}$ (2)
4 Experiments